INAF LBC+MODS queue observing, 2012 May 14-15

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: D. Thompson
Instrument: MODS + LBC

Summary

Clear skies at the beginning of the night.
At 5:30 Local Time we realised that LBC-RED was warming up (260K). Technical staff filled it up but it will take time to cool down. We NEED to start the night with the back-up MODS program.
At 09:20 Local Time LBC-RED just reached 187K and pressure 2.54E-05 ( Blue values: 182K and 1.8E-05). When we finish the MODS target we could switch to LBC if the seeing is good enough.
Humidity increased from 20% to 60% and seeing got bad (>2.00"). We stay with MODS until we have targets and then we switch to LBC.
Not a very productive night, because of the bad seeing.
The only good thing was that at the end of the night LBC was working and we did not have the elongation problems we had yesterday.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Turatto snNGC4490 red_grating 2hr15m seeing was variable between 1.0"-2.0" - see plot at the end of the log
Turatto sn2011aa dual_grating 2hr40m seeing was >2.0" - see plot at the end of the log
Standard Feige34 red/dual grating    
Standard Feige66 red/dual grating    
Piotto N6791 Vbess - i-sloan 50min seeing is blue 1.7"/red 1.45"

Details

02:51 acqMODS Feige34.acq mods1r 0003-0005

03:00 execMODS Feige34.obs mods1b 0003 / mods1r 0006-0007

03:15 acqMODS feige66.acq mods1r 0008-0011

03:26 execMODS feige66.obs mods1b 0004 / mods1r 0012-0013

03:36 acqMODS snNGC4490.acq mods1r 0014-0016

03:55 execMODS snNGC4490.obs mods1r 0017-0019
  • dimm seeing 1.6" / GCS 1.4"
  • GCS seeing 1.0" at the end of first exposure

05:29 dimm seeing 2.0" / GCS 2.0"

05:44 dimm seeing 2.4 / GCS 2.2"

06:13 GCS seeing ~2.2" near the end of the last exposure, so we decide to stay with MODS

06:20 Slew to sn2012aa.acq (Turatto)
  • acqMODS sn2012aa.acq
  • Acquisition fails! Guide star not found (out of range). We have to find another guide star

06:24 acqMODS sn2012aa.acq again mods1r 0020-0022
  • Problem with he guide star - it is unstable and jumpy because of bad seeing >2.0"
Send the offset to the Telescope (Y|N)?: Y
   Sending offset...
*** ERROR: offsetxy IIF error - OffsetPointing result status: Warning, guide star not stable in guide box
   Offset command returned error - /home/MODSeng/bin/modsAlign aborting

06:35 execMODS sn2012aa.obs mods1b 0005-0007 / mods1r 0023-0025
  • Seeing at the end of the 2nd exposure ~2.3"
  • dimm seeing 2.6"

08:44 Seeing is still >2.0". We repeat the exposures of this target until it reaches the PI limit for the airmass.
  • execMODS sn2012aa.obs mods1b 0008 / mods1r 0026

09:25 abort after the first exposure because we are at low elevation
  • we take a field image to see the position of the source mods1r 0027

09:35 No more targets for MODS - We switch to LBC

09:51 WD_N6791_FOCUS.ob

09:55 dofpia, /X2, /first

10:05 WD_N6791_LONG_05
  • seeing measured on images : blue 1.8" - red: 1.3"
  • dimm seeing 1.4"

10:37 dofpia at the filed - too many pupils, the fit is not correct
  • we slew to the focus field
  • dofpia

10:43 WD_N6791_LONG_06

11:13 Start with flats 11:35 SkyFlat_Vi_pa0 (scale 15) 11:41 SkyFlat_Vi_pa180 (scale 8). Aborted: too high 11:44 SkyFlat_Vi_pa180 (scale 4) 11:48 SkyFlat_Vi_pa0 (scale 1) 11:52 SkyFlat_Sr_pa0 (scale 2). Aborted: too high 11:54 SkyFlat_Sr_pa0 (scale 0.5)

12:00 Close dome - Start MODS calibrations for LS
  • execMODS grflats1.0.cal mods1b 0009-0018 / mods1r 0028-0042
  • execMODS grflats1.0red.cal mods1r 0043-0047

12:47 End of Night

  • 20120515.jpg:
    20120515.jpg

-- KonstantinaBoutsia - 15 May 2012
Topic revision: r8 - 26 May 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback